### v2

# Orbital decay candidates reconsidered: WASP-4 b is not decaying and Kepler-1658 b is not a planet

Abstract
The fate of hot Jupiters is thought to be engulfment by their host stars, the outcome of tidal orbital decay. Transit timing has revealed a few systems with apparently shrinking orbital periods, but such signals can be mimicked by light travel-time effects (LTTE) of a distant companion. Combining transit timings with precise radial-velocity data, including new data, we reassessed three reported cases of orbital decay: WASP-4, WASP-12, and Kepler-1658. For WASP-4, the period change is best explained by LTTE due to a $10~M_{\rm Jup}$ companion at 4 AU, with no need to invoke orbital decay. For WASP-12, in contrast, the data firmly exclude LTTE and confirm genuine orbital decay. For Kepler-1658, spectroscopic and photometric anomalies reveal the “planet” to be an eclipsing K/M binary bound to the F-type primary, with LTTE explaining the observed period change. Thus, among the known hot Jupiters, only WASP-12 b currently shows compelling evidence for orbital decay.
exoplanets: exoplanet systems, exoplanet dynamics, exoplanet astronomy
1 Introduction
Hot Jupiters have famously small orbits, some of which are only 3-4 times larger than their host stars. These hottest of the hot Jupiters are expected to be relatively short-lived on cosmological timescales, slowly spiraling into their host stars as tidal interactions sap their orbital energy and angular momentum. In close stellar binaries, tidal interactions usually lead to spin-orbit synchrony (Hut, 1980), but a hot Jupiter’s angular momentum is insufficient to spin a Sun-like star into synchrony (Rasio et al., 1996; Levrard et al., 2009), especially since magnetized winds steadily slow the star’s rotation (Barker & Ogilvie, 2009). However, the timescale of tidal orbital decay remains uncertain because it depends on poorly understood mechanisms of tidal dissipation within stars, such as inertial-wave damping and turbulent convection (see, e.g., Ogilvie, 2014).
Several population-level clues support the idea that tidal decay does occur. Sun-like stars with hot Jupiters rotate faster, on average, than comparable stars without such planets (Brown, 2014; Maxted et al., 2015; Penev et al., 2018; Tejada Arevalo et al., 2021), consistent with the tidal transfer of angular momentum. Kinematic age studies show that hot-Jupiter hosts are systematically younger than matched field stars (Hamer & Schlaufman, 2019; Miyazaki & Masuda, 2023), implying the planets are destroyed while their stars are still on the main sequence. More dramatically, the infrared transient reported by De et al. (2023) has been interpreted as the aftermath of a giant planet’s engulfment.
Direct evidence for orbital decay has come from transit timing. The best studied case is WASP-12 b, whose transit period is shrinking by about 30 ms/yr (Maciejewski et al., 2016; Patra et al., 2017; Yee et al., 2020). However, secular changes in the transit period can result from light travel-time effects (LTTE) of a distant massive companion. Also known as the “Rømer delay,” the effect is familiar in pulsar timing and triple-star systems (Edwards et al., 2006; Borkovits et al., 2015). When a star and its transiting planet are being pulled toward us, the observed transit period shrinks. Because radial-velocity (RV) measurements directly track the star’s line-of-sight motion, they provide a useful test of whether LTTE rather than orbital decay is responsible for a shrinking transit period.
Yee et al. (2020) used RV data to rule out LTTE as the cause of WASP-12 b’s shrinking period. Two other systems, WASP-4 b and Kepler-1658 b, also have apparently shrinking periods Although LTTE have been considered in both cases, and some authors have argued that they have been ruled out, no joint analysis of transit times and RVs has yet been presented.
In this paper, we report new transit-timing and RV observations of WASP-4 b and Kepler-1658 b, together with joint analyses designed to test the LTTE hypothesis. For illustrative purposes, we also revisit WASP-12 b, for which strong arguments have already been lodged against the LTTE hypothesis. The paper is organized as follows. Section 2 presents the relevant equations for the timing + RV analysis. Section 3 describes the case of WASP-4 b, and Section 4 provides the contrasting example of WASP-12 b. Section 5 describes our investigation of Kepler-1658 b and features a plot twist: we found evidence that the transits are not caused by a hot Jupiter, but rather by a low-mass eclipsing binary orbiting a brighter and more massive star. Brief reflections on the results are given in Section 6.
2 Mathematical preliminaries
The line-of-sight coordinate displacement of a star from the center of mass due to a single companion is
$z=\frac{a(1-e^{2})}{1+e\cos\theta}~\sin I~\sin(\theta+\omega),$ (1)
where $a$ is the semi-major axis, $e$ is the eccentricity, $I$ is the inclination, $\omega$ is the argument of periapse, and $\theta$ is the true anomaly. The time dependence of $\theta$ is governed by
$\tan\frac{\theta}{2}$ $=\sqrt{\frac{1+e}{1-e}}\tan\frac{E}{2}~~{\rm and}$ (2)
$E$- $e\sin E=n\left(t-t_{\rm p}\right),$ (3)
where $E$ is the eccentric anomaly, $t_{\rm p}$ is the time of periapse passage, and $n$ is the mean motion ($2\pi$ divided by the orbital period $P$). By taking the time derivative, one obtains the radial-velocity equation,
$\dot{z}=\frac{2\pi a}{P\sqrt{1-e^{2}}}\,\sin I\left[\cos(\theta+\omega)+e\cos\omega\right].$ (4)
A few changes in parameterization are useful. First, we rewrite $a\sin I$ in terms of the radial-velocity semi-amplitude,
$K\equiv\frac{2\pi a}{P\sqrt{1-e^{2}}}\,\sin I,$ (5)
because $K$ is more directly observable. Second, instead of using $t_{\rm p}$ to pinpoint the star’s location at a particular time, we use the true anomaly at inferior conjunction,
$\theta_{\rm c}\equiv\frac{\pi}{2}-\omega,$ (6)
because it allows a more graceful approach to the $e\rightarrow 0$ limit in which $t_{\rm p}$ is ill-defined. With these changes, $z$ and $\dot{z}$ can be written
$z$ $=$ $\frac{K}{n}\left(1-e^{2}\right)^{\!3/2}\,\frac{\cos(\theta-\theta_{\rm c})}{1+e\cos\theta},$ (7)
$\dot{z}$ $=$ $K\left[e\sin\theta_{\rm c}-\sin(\theta-\theta_{\rm c})\right].$ (8)
Assuming the star has a hot Jupiter and a wider-orbiting companion with negligible mutual interactions, the observed radial velocity is
$v=v_{0}+\gamma+\dot{z}_{1}(t)+\dot{z}_{2}(t),$ (9)
where $v_{0}$ is the radial velocity of the planetary system’s center of mass relative to the Solar System, $\gamma$ is a spectrograph-dependent additive offset [^1] The $\gamma$ parameter is necessary because spectrographs measure radial-velocity variations at the m/s level, but the accuracy of the absolute radial velocity scale is usually no better than 100 m/s, leading to systematic offsets between different spectrographs., and $\dot{z}_{1}$ and $\dot{z}_{2}$ are the contributions from the hot Jupiter and wider-orbiting companion, respectively.
In the same scenario, the transit times are
$t_{k}=t_{0}+kP_{1}+\frac{z_{2}(t_{0}+kP_{1})-z_{2}(t_{0})}{c},$ (10)
where $t_{0}$ is the transit time of an arbitrarily chosen orbit, $k$ is an integer, and the last term is the LTTE correction.
It will be useful to simplify these equations for the case when the period of the wide-orbiting companion is longer than the time span of the data. We designate $t=0$ as the time of the first data point and perform a Taylor expansion in the small parameter $2\pi t/P_{2}=n_{2}t$:
$\dot{z}_{2}=\sum_{j=0}^{\infty}\beta_{j}(n_{2}t)^{j}.$ (11)
For the special case $e=0$, the first four coefficients are
$\beta_{0}=$ $-K_{2}\sin\phi_{0}$ $~{\rm(velocity),}$ (12)
$\beta_{1}=$ $-K_{2}n_{2}\cos\phi_{0}$ $~{\rm(acceleration),}$ (13)
$\textstyle\beta_{2}=$ $\tfrac{1}{2}K_{2}n_{2}^{2}\sin\phi_{0}$ $~{\rm(jerk),\,and}$ (14)
$\beta_{3}=$ $\tfrac{1}{6}K_{2}n_{2}^{3}\cos\phi_{0}$ $~{\rm(snap),}$ (15)
where $\phi_{0}\equiv\theta(0)-\theta_{\rm c}$. The corresponding approximation for the transit times is obtained using Equation 10 and the time integral of Equation 11:
$t_{k}\approx t_{0}+kP_{1}+\frac{1}{n_{2}c}\sum_{j=1}^{\infty}\frac{k^{j}}{j}\beta_{j-1}(n_{2}P_{1})^{j}.$ (16)
3 WASP-4
[…]
Whether or not orbital decay has been detected for WASP-4 b has been the subject of debate. Early timing studies showed the transit period to be shrinking at a rate of order $\delta P/P=10~{\rm ms/yr}$ (Bouma et al., 2019; Southworth et al., 2019), although the robustness of this finding was questioned (Baluev et al., 2019). Additional RV data suggested the star is accelerating a rate compatible with the LTTE hypothesis (Bouma et al., 2020), but concerns were raised about systematic errors (Baluev et al., 2020). With additional data, Turner et al. (2022) confirmed a shrinking period and argued that the LTTE hypothesis was firmly ruled out, but an error in their analysis was later identified, showing that LTTE was still viable (Harre & Smith, 2023). Most recently, Baştürk et al. (2025) reported further transit times and concluded that orbital decay is the best explanation.
This sequence of claims and counter-claims illustrates the difficulty of detecting slight period changes and disentangling orbital decay from LTTE. We suggest that an underlying reason for the confusion is that none of the prior studies performed a joint fit of the radial-velocity and transit-timing data.
3.1 Radial Velocities
Radial velocities spanning 13 years are available from three spectrographs: CORALIE [^2] CORALIE is mounted on the Euler 1.2 m telescope at La Silla Observatory, Chile. (Wilson et al., 2008; Baluev et al., 2019), HARPS [^3] High Accuracy Radial velocity Planet Searcher (HARPS) is mounted on the La Silla 3.6 m telescope.(Triaud et al., 2010; Baluev et al., 2019), and HIRES [^4] The High Resolution Echelle Spectrometer (HIRES) was used on the Keck I 10 m telescope on Mauna Kea, Hawaii. (Bouma et al., 2020). We began with the compilation provided by Turner et al. (2022) and applied the following filters to construct a clean dataset:
  • We excluded all velocities measured during transits, to avoid modeling the Rossiter-McLaughlin effect.
  • We omitted a single CORALIE velocity (BJD 2456149.87585) with an unusually large uncertainty (59 m/s, compared to the typical 15–25 m/s).
  • We added a new HIRES velocity from December 26, 2020 (BJD 2459189.71567), extending the HIRES time span by 1.3 years.
  • To mitigate the effects of correlated errors due to stellar rotation (Husnoo et al., 2012), we “thinned” the data by retaining only one point every five days.
The final dataset comprised 37 RVs: 23 from CORALIE, 5 from HARPS, and 9 from HIRES. The data are displayed in the top left panel of Figure 1. The peak-to-peak variations of about 400 m/s are caused by the hot Jupiter.
3.2 Transit Times
WASP-4 was observed in Sectors 2, 28, 29, 69, and 96 of NASA’s Transiting Exoplanet Survey Satellite (TESS) mission (Ricker et al., 2015), and all the data are available with 120 s time sampling. The Sector 96 data have not previously been analyzed in the literature. We used light curves processed by the TESS Science Processing and Operations Center (SPOC) to measure transit times with the procedures of Ivshina & Winn (2022), obtaining 70 transit times with typical uncertainties of 25 s.
Many other transit times have been reported for WASP-4 b. A challenge of long-term timing analyses is deciding which measurements to include and how to assign realistic error bars, given the variable quality of ground-based data, the use of heterogeneous instruments and methods, and occasional misunderstandings about timing systems and transcription errors (see, e.g., Adams et al., 2024, for a careful examination of several such cases). Our approach was to
  • include the pre-TESS transit times compiled by Ivshina & Winn (2022), which are drawn exclusively from the peer-reviewed literature,
  • exclude the ground-based data obtained after the first TESS observation,
  • discard the three data points with formal uncertainties $\geq$60 s, and
  • impose a minimum uncertainty of 25 s when fitting the data, to prevent any single data point from having undue weight.
After applying these criteria, the pre-TESS dataset comprises 66 transit times. The lower left panel of Figure 1 shows the deviations between the measured transit times and the best-fit constant-period model, revealing the downward-curving trend that has been interpreted as evidence for orbital decay.
Figure 1: WASP-4. Upper left.—Radial velocities. Lower left.—Deviations between transit times and the best-fit constant-period model ($P=1.338231293$ days, $t_{\rm c}=2456180.558583$ BJDTDB). In this panel and others, the black curve is the best-fit model including the hot Jupiter and a wide-orbiting companion, and the gray curves are random draws from the posterior. Upper right.—Radial velocities as a function of folded time (i.e. relative to the nearest transit time), after subtracting the best-fit contribution from the wide-orbiting companion. Lower right.—Radial velocities after subtracting the best-fit contribution of the hot Jupiter.
3.3 Timing+RV analysis
Previous work established that WASP-4 b’s orbit is nearly circular (Husnoo et al., 2012), as expected from tidal interactions, and that the radial-velocity trend evolves gradually over the 13-year baseline (Bouma et al., 2020). With this in mind, we adopted a model consisting of a circular orbit for the hot Jupiter, and a low-order polynomial approximation for the motion induced by the distant companion:
$\dot{z_{1}}$ $=-K_{1}\sin\left[\frac{2\pi}{P_{1}}(t-t_{\rm c,1})\right]~{\rm and}$ (17)
$\dot{z}_{2}$ $=\beta_{1}t+\beta_{2}t^{2}+\beta_{3}t^{3}.$ (18)
The model had nine free parameters: $K_{1}$, $P_{1}$, $t_{\rm c,1}$, the three $\beta$ parameters, and the three $\gamma$ offsets (one per spectrograph). The best-fit yielded $\chi^{2}=158$ with 164 degrees of freedom, indicating an excellent fit. Posterior sampling was performed using the Markov Chain Monte Carlo (MCMC) algorithm of Goodman & Weare (2010) as implemented in emcee (Foreman-Mackey et al., 2013). To allow for possible underestimation of uncertainties, we introduced a velocity jitter term for each spectrograph. Table 1 gives the results and Figure 1 shows the corresponding fits. The model provides a good fit to all the observables.
Table 1: WASP-4 model parameters
Parameter Best-fit Marginalized Units
value posterior
$P_{1}$ 1.33823140 $1.338231397^{+0.000000024}_{-0.000000024}$ days
[…]
$K_{1}$ 237.8 $237.8^{+2.6}_{-2.7}$ m/s
$\beta_{1}$ 16.9-16.9 $16.4^{+1.4}_{-1.4}$ m/s/year
$\beta_{2}$ 4.7-4.7 $4.56^{+0.65}_{-0.66}$ m/s/year2
$\beta_{3}$ $6.5$ $5.3^{+1.6}_{-1.7}$ m/s/year3
$\gamma_{\rm CORALIE}$ 29.1-29.1 $34.0^{+6.4}_{-6.4}$ m/s
$\gamma_{\rm HARPS}$ 60.6-60.6 $62.0^{+7.5}_{-7.9}$ m/s
$\gamma_{\rm HIRES}$ 44.9-44.9 $44.4^{+3.8}_{-3.8}$ m/s
$\sigma_{\rm CORALIE}$ $6.5$ $9.4^{+5.7}_{-5.9}$ m/s
$\sigma_{\rm HARPS}$ $7.3$ $10.0^{+5.2}_{-3.9}$ m/s
$\sigma_{\rm HIRES}$ $2.2$ $5.4^{+3.0}_{-2.1}$ m/s
$K_{2}$ $23.0$ $22.5^{+7.9}_{-2.4}$ m/s
$P_{2}$ $7.2$ $7.10^{+2.03}_{-0.74}$ years
$a_{2}$ $3.6$ $3.56^{+0.60}_{-0.23}$ AU
$m_{2}\sin I_{2}$ $9.4$ $9.2^{+3.8}_{-1.2}$ $M_{\rm Jup}$
Note. — Parameters below the line were constructed from parameters above the line assuming $e=0$. Data in Column 3 are based on the 16th, 50th, and 84th percentiles of the cumulative posterior probability function for each parameter, marginalized over all other parameters.
The five parameters of the companion’s radial-velocity orbit ($K_{2}$, $P_{2}$, $e_{2}$, $\omega_{2}$, and $t_{\rm c,2}$) cannot be uniquely determined because the data provide only three constraints ($\beta_{1}$, $\beta_{2}$, and $\beta_{3}$). Nevertheless, we can estimate the companion’s minimum mass and orbital separation by assuming $e_{2}\approx 0$ and using Equations 1215 to find
$n_{2}=\sqrt{-\frac{2\beta_{3}}{\beta_{1}}}~~{\rm and}~~K_{2}=\frac{\beta_{1}}{n_{2}}\sqrt{1+\left(\frac{2\beta_{2}}{n_{2}^{2}}\right)^{\!\!2}}.$ (19)
Adopting a stellar mass of $0.9\,M_{\odot}$ (Wilson et al., 2008), we infer a companion with minimum mass $\sim$10 $M_{\rm Jup}$ at $\sim$4 AU (Table 1, Figure 2). The posterior distributions for $a_{2}$ and $m_{2}\sin i_{2}$ have long tails extending toward larger masses and wider separations, but the radial acceleration of the primary star is constrained more tightly:
$\frac{m_{2}\sin I_{2}}{a_{2}^{2}}=0.704_{-0.029}^{+0.068}~\frac{M_{\rm Jup}}{{\rm AU}^{2}}.$ (20)
Allowing the companion to have a nonzero eccentricity would modify these estimates by factors of order $e_{2}$.
In summary, a super-Jupiter or brown dwarf companion at several AU naturally explains the transit-timing and RV variations of WASP-4, without invoking orbital decay.
Figure 2: Constraints on the wide-orbiting companion of WASP-4 b. Shown are the posteriors for minimum mass and orbital separation, assuming the orbit is nearly circular. The contours enclose 68%, 95%, and 99.7% of the probability.
4 WASP-12
Figure 3: WASP-12. Upper left.—Radial velocities. Lower left.—Deviations between transit times and the best-fit constant-period model ($P=1.091418667$ days, $t_{\rm c}=2456305.45552$ BJDTDB). In this panel and others, the black curve is the best-fit model including the effects of the hot Jupiter and a wide-orbiting companion, and the gray curves are random draws from the posterior. Upper right.—Radial velocities as a function of folded time (i.e. relative to the nearest transit time), after subtracting the best-fit contribution of the wide-orbiting companion. Lower right.—Radial velocities after subtracting the best-fit contribution of the hot Jupiter.
WASP-12 has been classified as 6500 K main-sequence star with mass 1.4 $M_{\odot}$ and 1.7 $R_{\odot}$ (Collins et al., 2017; Bailey & Goodman, 2019), although there are suggestions it may be somewhat less massive and more evolved (Weinberg et al., 2017). The hot Jupiter WASP-12 b, discovered by Hebb et al. (2009), has mass 1.5 $M_{\rm Jup}$, radius $1.9~R_{\rm Jup}$, and period 1.1 days. With an orbital separation only three times the stellar radius, WASP-12 b lies precariously close to the Roche limit (Antonetti & Goodman, 2022). We do not attempt to review the vast WASP-12 literature (see Haswell, 2018, for an overview). Our aim was to apply the same joint timing + RV analysis used for WASP-4 to test the viability of the LTTE hypothesis for this system.
TESS data with 120 s sampling are available from Sectors 20, 43, 44, 45, 71, and 72. We re-analyzed the SPOC light curves and combined the measured transit times with the pre-TESS compilation of Ivshina & Winn (2022). As before, we excluded ground-based data obtained after the first TESS observation, imposed a minimum uncertainty of 25 s and discarded times with formal uncertainties exceeding 60 s.
There are three prior sources of radial velocities: the SOPHIE [^5] SOPHIE is the Spectrographe pour l’Observation des Phénomènes des Intérieurs stellaires et des Exoplanètes, mounted on the 1.9 m telescope at Haute-Provence Observatory, France. data of Hebb et al. (2009), the HARPS-N [^6] HARPS-N is the High Accuracy Radial velocity Planet Searcher for the Northern hemisphere, fed by the 3.6 m Telescopio Nazionale Galileo at Roque de los Muchachos Observatory on the La Palma, Canary Islands, Spain. data of Bonomo et al. (2017), and the HIRES data of Yee et al. (2020). We supplemented these with 20 new velocities obtained with the NEID [^7] NEID stands for NN-EXPLORE Exoplanet Investigations with Doppler Spectroscopy. It is a stabilized optical spectrograph with resolution $\approx$120,000 (Schwab et al., 2016) mounted on the WIYN 3.5 m telescope at Kitt Peak, Arizona. spectrograph in 2024 and 2025. The NEID spectra were processed with version 1.4.2 of the data reduction pipeline. [^8] https://neid.ipac.caltech.edu/docs/NEID-DRP/ Spectra obtained during transits were excluded, and the time series was thinned to enforce a five-day minimum time spacing.
Figure 3 shows the RVs and transit timing deviations. The 400 m/s RV variations are due to the hot Jupiter, with no obvious long-term trend. The transit timings, however, display a clear quadratic trend.
To test the LTTE hypothesis, we used a model similar to that used for WASP-4, modified only to account for the RV anomalies caused by the star’s tidal distortion (Arras et al., 2012). For small distortions, the effect introduces a sinusoidal signal with a period of $P_{1}/2$, an amplitude of a few m/s, and a phase shift of 90 relative to the main signal. We modeled this effect by the simple expedient of assigning the planet a fictitious eccentricity with $\omega=-\pi/2$ (for more details, see Arras et al., 2012; Maciejewski et al., 2020).
The best-fit model is shown in Figure 3 (dashed black curves). The fit is very poor, with with $\chi^{2}=1{,}531$ and 323 degrees of freedom. The model attempts to reproduce the quadratic timing trend with a suitably strong line-of-sight acceleration (lower left) but is stymied by the constraint imposed by contemporaneous RV data (lower right), especially the HIRES data (green points).
In short, the joint timing + RV analysis firmly rules out LTTE as the explanation for the shrinking transit period of WASP-12 b. This result corroborates earlier studies and provides a useful contrast with WASP-4 b, where LTTE offers a compelling explanation.
5 Kepler-1658
Figure 4: Kepler-1658. Upper left.—Radial velocities. Lower left.—Deviations between transit times and the best-fit constant-period model ($P=3.849363145$ days, $t_{\rm c}=2455005.925555$ BJDTDB). In this panel and others, the black curve is the best-fit model including the effects of the hot Jupiter and a wide-orbiting companion, and the gray curves are random draws from the posterior. Upper right.—Radial velocities as a function of folded time (i.e. relative to the nearest transit time) after subtracting the best-fit contribution of the wide-orbiting companion. Lower right.—Radial velocities after subtracting the best-fit contribution of the hot Jupiter.
Kepler-1658 was first described in detail by Chontos et al. (2019), using data from NASA’s Kepler mission (Borucki et al., 2010) and follow-up observations. Early in the mission, a sequence of transits was detected with a period of 3.9 days and a depth of 0.13%. The source was inconsistently classified in mission catalogs, appearing as a planet candidate in some and a false positive in others. Confirmation of Kepler-1658 b as a planet was based mainly on the detection of radial-velocity variations with the expected period and phase and an amplitude consistent with a planetary mass ($\approx$ 6 $M_{\rm Jup}$). The host star is a subgiant with effective temperature 6200 K, mass $1.4\,M_{\odot}$, and radius $2.9\,R_{\odot}$. Because tidal dissipation rates are expected to increase as a star evolves into the subgiant phase, the system was highlighted as a favorable target for detecting orbital decay.a
Vissapragada et al. (2022) analyzed TESS data from Sectors 41, 54, and 55, as well as two ground-based observations with the Wide Field Infrared Camera (WIRC) on the Hale 5 m telescope at Palomar Observatory, California. When combined with the Kepler data, these transit times revealed apparent orbital decay at a rate of 130 ms/yr, five times faster than WASP-12 b and greatly exceeding theoretical expectations (Barker et al., 2024).
Could the decreasing transit period be caused by the light travel-time effect of an outer companion? Vissapragada et al. (2022) argued that the existing RV data were sufficient to rule out this possibility. Our joint timing + RV analysis, described below, shows that this conclusion was premature. More fundamentally, our analysis of new data and re-analysis of older data point to a different interpretation of Kepler-1658 altogether: it is not a hot Jupiter system, but rather an unresolved triple star system.
5.1 Radial Velocities
Chontos et al. (2019) used TRES [^9] TRES is the Tillinghast Reflector Echelle Spectrograph, mounted on the 1.5 m telescope at Mt. Hopkins, Arizona. to measure 23 RVs with typical uncertainties of 150 m/s. Three measurements were made in 2009 and 2010, and the other 20 came after a seven-year gap. To this dataset we added 22 velocities from NEID obtained in 2024–2025. The official NEID pipeline reduction (v1.4.3) produced velocities with anomalously large uncertainties and no sign of the known 3.9-day periodicity. [^10] In retrospect, this might have been due to the spectral anomalies discussed in Section 5.4. We therefore performed a custom extraction using the SERVAL template-matching code (Zechmeister et al., 2018), as adapted for NEID by Stefànsson et al. (2022). The resulting velocities had uncertainties of 50 m/s, more in line with expectations.
The RV data are shown in the upper left panel of Figure 4. The peak-to-peak variations of about 1,000 m/s have been attributed to the hot Jupiter. Beyond this, as noted by Vissapragada et al. (2022), there is no evidence for a long-term trend. The constraint is weaker than it may appear, though, because the TRES and NEID data can be shifted relative to each another by an arbitrary constant. The primary leverage on any RV trend comes from the consistency between the earliest three TRES velocities and those obtained seven years later.
5.2 Transit Times
We re-analyzed the TESS light curves from Sectors 41, 54, and 55, and incorporated newly TESS data from Sectors 74, 75, 81, and 82. Because single transits are barely detectable in the TESS photometry, we abandoned attempts to measure individual transit times. Instead, we grouped the data into four multi-transit blocks: Sector 41, Sectors 54 & 55, Sectors 74 & 75, and Sectors 81 & 82. Within each block, the light curve was fitted assuming strict periodicity, with the key parameter being the transit time of the event closest to the block center. There was no requirement for strict periodicity across different blocks. This approach offers greater robustness at the expense of sensitivity to short-term variations. Vissapragada et al. (2022) used a similar procedure to analyze the Kepler data, treating the observing quarters as independent blocks. We adopted without modification their reported transit times from Kepler and from WIRC.
The timing deviations are displayed in the lower left panel of Figure 4. Even without reference to the fitted curves, the data indicate that the average transit period during the Kepler mission was longer than during the TESS era. This was the basis of the claim of orbital decay.
5.3 Analysis
The case of Kepler-1658 differs from that of WASP-4 in two important respects: no long-term RV trend is evident, and the datasets are relatively sparse and irregularly spaced. The long observational gaps leave open the possibility of variations that are not well approximated by low-order polynomial functions. For this reason, we modeled the effects of the hypothetical companion with a full Keplerian orbit parameterized by $\{K_{2},P_{2},t_{\rm c,2},e_{2},\omega_{2}\}$, resulting in 10 free parameters and 62 data points.
We scanned $10^{4}$ trial values of $P_{2}$, log-uniformly distributed between $7$ and $10^{5}$ days. For each choice of $P_{2}$, we optimized the other parameters and recorded the value of $\chi^{2}$. The optimal solution was found at $P_{2}\approx 2500$ days, with $\chi^{2}=187$ with 52 degrees of freedom — a poor fit. We proceeded under the tentative assumption that the RV uncertainties had been underestimated, which seemed plausible since precise RV extraction codes are not optimized for broad-lined stars such as Kepler-1658 ($v\sin i\approx 34$ km/s). By adding terms in quadrature of $\sigma_{\rm TRES}=180$ m/s and $\sigma_{\rm NEID}=100$ m/s, the $\chi^{2}$ of the best-fit model was lowered to equal the number of degrees of freedom.
[…]
While this exercise demonstrated that the LTTE hypothesis cannot be ruled out as the explanation for the observed change in the transit period, the models are unconvincing. The required jitter terms are large, and the models seem to “cheat” by arranging for the strongest variations to occur during gaps in the dataset. This prompted us to re-examine the Kepler-1658 data more closely.
5.4 Problems with the planet hypothesis
We began by checking for consistency of the RV semi-amplitudes measured by the two different instruments. Fitting each dataset with a sinusoidal function having the known period and phase [^11] When fitting the TRES data, we did not include the three earliest data points from 2009, to reduce the possible effects of any long-term trend. yielded $K_{\rm TRES}=561\pm 75$ m/s and $K_{\rm NEID}=318\pm 40$ m/s, a significant difference of $\Delta K=243\pm 85$ m/s. The discrepancy can be seen in Figure 4, where the blue and orange points trace different amplitudes.
We then inspected the NEID cross-correlation functions (CCFs), obtained by correlating each spectrum against an F-star template. The CCFs exhibit night-to-night variations at the percent level (Figure 5). This brought to mind a case presented by Mandushev et al. (2005), in which a rapidly rotating F star was initially thought to have a transiting brown dwarf, but the signal was eventually shown to arise from the blended light of an F star and a fainter eclipsing binary. The shifting lines of the binary produced subtle distortions within the F star’s broad lines that were mistaken for RV variations.
Suspecting that Kepler-1658 could be a Mandushev-type false positive, we looked for evidence of orbital motion in the shifting pattern of CCF residuals. We subtracted the median NEID CCF from each individual CCF, and then cross-correlated the residuals. The resulting velocity shifts, shown in the right panel of Figure 5 follow a sinusoid with the same period and phase as the transit signal, and a semi-amplitude of about 20 km/s. This is precisely the signature expected of a blended eclipsing binary masquerading as a planetary companion.
Figure 5: Spectral line variations of Kepler-1658. Left.—Representative cross-correlation functions (CCFs) based on NEID spectra. They are centered on the systemic velocity of 24-24 km/s and have widths of about 30 km/s due to stellar rotation. There are also night-to-night shape variations. Middle.—Isolation of CCF variations. The median NEID CCF was subtracted from individual CCFs, revealing anomalies of amplitude $\sim$$10^{-2}$ that shift in velocity. Right.—Velocity shift of the pattern of residuals relative to an arbitrarily chosen night, as a function of time folded with the 3.8-day transit period. The variation is consistent with 20 km/s orbital motion of an unresolved binary with the expected period and phase.
Seeking corroborating evidence, we re-examined the Kepler photometry. A median light curve was constructed from all the 1-min data, phase-folded using the transit period, and phase-averaged to 2.3-min cadence (124 data points) to speed up further analysis. The standard deviation outside of transits was 22 parts per million (ppm), which we adopted as the per-point uncertainty. Fitting the light curve with the Mandel & Agol (2002) transit model gave $\chi^{2}=134.5$ with 117 degrees of freedom. While this is statistically acceptable, there are structured patterns in the residuals, especially during ingress and egress (Figure 6).
Motivated by the blended binary hypothesis, we fitted the light curve with a diluted model:
$F(t)=(1-\epsilon)+\epsilon\,\delta(t),$ (21)
where $\delta(t)$ is the Mandel & Agol (2002) model and $\epsilon$ is the fractional light contribution of the binary. Adding the single parameter $\epsilon$ lowered $\chi^{2}$ by 20.3 units and reduced the structure in the residuals, yielding $\epsilon=0.0173^{+0.0020}_{-0.0014}$.
Figure 7 shows that smaller and larger values of $\epsilon$ are disfavored. When $\epsilon$ is increased from the best-fit value of 0.0173, the model preserves the transit depth by reducing the radius ratio $r/R$, and preserves the transit duration by increasing the impact parameter $b$, but the fit to the subtle details of ingress and egress becomes worse. Lowering $\epsilon$ below the best-fit value fails because the transit duration cannot be preserved; the impact parameter is already at its minimum value.
Figure 6: Evidence for dilution in the light curve of Kepler-1658. Top.—Kepler light curve, with its native 1-minute sampling as well as the phase-averaged version used for model fitting. Bottom.—Residuals between the data and the best-fit transiting-planet model. The pattern of residuals is well-fitted by a model in which the “transits” are eclipses of a faint binary that contributes only 1.7% of the total light.
Figure 7: Constraints on dilution from the Kepler-1658 light curve. Top.—Goodness of fit, $\chi^{2}$, as a function of $\epsilon$, the fraction of the light contributed by the blended binary. Middle and bottom.—Corresponding values of the radius ratio between the two stars of the eclipsing binary, and the eclipse impact parameter. The data favor $\epsilon\approx 0.017$, $r/R\approx 0.27$, and $b\approx 0$.
In summary, the evidence against the planetary hypothesis for Kepler-1658 b seems compelling: inconsistent RV amplitudes between instruments, sinusoidal CCF variations with the same period and phase as the transits, and evidence of strong dilution in the Kepler light curve. Taken together, these findings point to a hierarchical triple system rather than a hot Jupiter.
5.5 Hierarchical triple-star model
Our next task was to identify plausible parameters for a hierarchical triple that reproduces all the relevant observations. We denote the broad-lined F star as component A, and the members of the 3.8-day eclipsing binary as B and C. The goal was to determine $M_{\rm B}$, $M_{\rm C}$, and the separation of A from the B/C pair, subject to five observational constraints:
  • the apparent transit and occultation depths observed by Kepler,
  • the $J$-band apparent transit depth seen in the WIRC light curve,
  • the dilution factor of $\sim$$10^{2}$ inferred from the Kepler light curve,
  • the $\sim$20 km/s sky-projected orbital speed indicated by the CCF residuals, and
  • the apparent decline of the eclipse period by 130 ms/yr.
To relate the stellar parameters to these observables, we assumed B and C are main-sequence dwarfs and employed the “Mamajek” zero-age main-sequence relations [^12] Table 6 of Pecaut & Mamajek (2013), as updated online at https://www.pas.rochester.edu/~emamajek/EEM_dwarf_UBVIJHK_colors_Teff.txt linking mass, radius, and absolute magnitudes in various bandpasses. We also needed transformations into Kepler magnitudes, which we obtained from Brown et al. (2011).
The depth of the secondary eclipse specifies the Kepler-band flux ratio between C and A:
$\delta_{\rm occ}=\frac{F_{\rm C}}{F_{\rm A}+F_{\rm B}+F_{\rm C}}\approx\frac{F_{\rm C}}{F_{\rm A}}.$ (22)
Since A’s properties are observed directly, we can use the flux ratio to calculate C’s Kepler-band absolute magnitude and the corresponding stellar mass. To match the observed depth of 61 ppm, star C should be an M dwarf with $M_{\rm Kep}=12.47$, $M_{J}=8.67$, mass $0.19\,M_{\odot}$, and radius $0.23\,R_{\odot}$.
The predicted rate of change of the period due to LTTE is
$\frac{\delta P}{P}\sim\frac{P}{c}\,\frac{GM_{\rm A}}{a^{2}}\sim 130~{\rm ms/yr}\left(\frac{M_{\rm A}}{1.4\,M_{\odot}}\right)\left(\frac{a}{50~{\rm AU}}\right)^{\!-2}\!\!\!.$ (23)
Consistency with the observed rate is achieved if the eclipsing binary is located $a\sim 50$ AU from the F star. At the Gaia-determined distance of 788 parsecs (Gaia Collaboration et al., 2023) [^13] https://gaia.ari.uni-heidelberg.de/singlesource.html the implied angular separation is $0.06$ arcsec, consistent with being unresolved.
The apparent transit depth is
$\delta_{\rm tra}\approx\frac{F_{\rm B}\left(\frac{R_{\rm C}}{R_{\rm B}}\right)^{\!2}}{F_{\rm A}+F_{\rm B}+F_{\rm C}},$ (24)
which can be considered a function of the single parameter $M_{\rm B}$ using the Mamajek relations. Figure 8 shows that $M_{\rm B}\approx 0.75\,M_{\odot}$ (a K dwarf) simultaneously satisfies the constraints on the Kepler-band transit depth (top) and the 1.7% flux contribution of the binary inferred from the diluted model of the Kepler light curve (bottom). With star B at $0.75\,M_{\odot}$, the orbital velocity is predicted to be 27 km/s, consistent with the sky-projected velocity of 20 km/s inferred from the CCF residuals.
The only tension is with the $J$-band transit depth, which is predicted to be almost twice as large as displayed in Figure 2 of Vissapragada et al. (2022). It should be noted, though, that the $J$-band light curve was constructed with a flexible model for systematics and a prior constraint enforcing the transit depth to be similar to the Kepler-band value. [^14] According to Vissapragada et al. (2022), the time series of WIRC aperture fluxes was fitted with a model having free parameters for the weights of 9-10 comparison stars and the strengths of correlations with instrumental parameters such as focus and centroid position. It seems plausible that allowing the $J$-band transit depth to vary freely would yield a broader range of values consistent with our prediction.
Figure 8: Kepler-1658 as a hierarchical triple star. The dependence of various observables is shown as a function of the mass of star B. Choosing $M_{B}\approx 0.75\,M_{\odot}$ (light gray band) is compatible with all the observables except the $J$-band apparent transit depth (see the text for details).
The triple-star model also provides natural explanations for a few peculiarities of the planet hypothesis:
  • For a transiting planet, the bulk density of the star can be calculated as (Seager & Mallén-Ornelas, 2003; Winn, 2010)
    $\rho_{\star}=\frac{3\pi}{GP^{2}}\left(\frac{a}{R}\right)^{\!3}.$ (25)
    Fitting the Kepler light curve with $\epsilon=0$ gives $\rho_{\star}=0.19\pm 0.04$ g/cm3, which disagrees by 4.5-$\sigma$ with the value of $0.083\pm 0.008$ g/cm3 determined by Chontos et al. (2019) using the independent technique of asteroseismology. [^15] By allowing the hot Jupiter’s orbit to be eccentric, Chontos et al. (2019) relieved but did not eliminate this tension; in their joint model the median of the posterior for $\rho_{\star}$ differs by 3-$\sigma$ from the asteroseismic value (cf. their Tables 2 and 4). In the triple-star model, the eclipse parameters are unrelated to the bulk density of star A and there is no tension.
  • If interpreted as reflected light from the hot Jupiter, the secondary eclipse depth of 62 ppm implies an albedo of $0.724_{-0.081}^{+0.090}$. While this is physically possible, most other albedo measurements are $\lesssim 0.3$ (see, e.g., Figure 10 of Wong et al. 2021). In the triple-star model, there is nothing unusual about the secondary eclipse depth.
  • The Kepler light curve shows 0.1% variations with a period of 5.7 days. For this to be due to rotation of the F star, it is necessary to invoke high-latitude starspots and 20-40% latitudinal differential rotation (Chontos et al., 2019). In the triple-star model, the modulation can be attributed to star B.
6 Discussion
Long-term transit timing of hot Jupiters offers many possible rewards: the detection of orbital decay, the discovery of additional orbiting bodies (see, e.g., Korth et al., 2024; Yang et al., 2025), and possibly even the effects of the planet’s own tidal deformation (Ragozzine & Wolf, 2009). Another attraction of transit timing is that relatively small telescopes – even amateur telescopes – can make scientifically valuable contributions. The arrival of high-quality all-sky data from the TESS mission has improved the reliability and expanded the accessibility of transit-timing studies. However, the cases examined in this paper underscore that transit timing by itself is usually not enough to earn the scientific rewards. Light travel-time effects from outer companions can mimic orbital decay, and long-term RV monitoring is essential to distinguish these possibilities.
A more specific lesson from the work described in this paper is that measuring or placing an upper bound on a linear RV trend is not sufficient to rule out the LTTE hypothesis. For WASP-4, this was because the RV trend was not well described as linear over the relevant time span. For datasets with sparse coverage, similar to that of Kepler-1658, large gaps admit the possibility of shorter-term variations that are poorly described by linear trends.
Finally, the Kepler-1658 story is a reminder that false positives remain a real hazard, even after two decades of experience with wide-field transit surveys. Blended eclipsing binaries bound to bright primaries, especially when the primaries are rapid rotators, can evade standard vetting procedures and mimic planetary signals to high precision. Careful scrutiny of spectral line shapes and light-curve details is essential.
Of the three systems considered here, only WASP-12 b shows persuasive evidence for orbital decay, supported not only by joint timing + RV analyses but also the observation that the occultation period is shrinking in step with the transit period (Yee et al., 2020). As the timing baseline from TESS grows longer, and after the PLATO and Roman missions commence (Rauer et al., 2025; Carden et al., 2025) additional candidates for orbital decay are likely to emerge. Interpreting them will require persistence, long-term RV monitoring, and careful vetting to distinguish genuine orbital changes from confounding effects.
7 Acknowledgments
J.N.W. gratefully acknowledges support from the NASA TESS project and a NASA Keck PI award. The work described in this paper was based in part on observations with the NEID instrument on the WIYN 3.5m telescope at Kitt Peak National Observatory. Kitt Peak is a facility of NSF’s NOIRLab, managed by the Association of Universities for Research in Astronomy (AURA). The WIYN telescope is a joint facility of NOIRlab, Indiana University, the University of Wisconsin-Madison, Pennsylvania State University, Purdue University, and Princeton University. NEID was funded by the NASA-NSF Exoplanet Observational Research (NN-EXPLORE) partnership and built by Pennsylvania State University. The NEID archive is operated by the NASA Exoplanet Science Institute at the California Institute of Technology. NN-EXPLORE is managed by the Jet Propulsion Laboratory, California Institute of Technology under contract with the National Aeronautics and Space Administration. The authors thank the NEID Queue Observers and WIYN Observing Associates, as well as the California Planet Search team supporting precise RV acquisition with the Keck I telescope, for their dedicated and skillful execution of the observations.